### v1

# CLARITY: Contextual Linguistic Adaptation and Accent Retrieval for Dual-Bias Mitigation in Text-to-Speech Generation

Abstract
Instruction-guided text-to-speech (TTS) research has reached a maturity level where excellent speech generation quality is possible on demand, yet two coupled biases persist: accent bias, where models default to dominant phonetic patterns, and linguistic bias, where dialect-specific lexical and cultural cues are ignored. These biases are interdependent, as authentic accent generation requires both accent fidelity and localized text. We present CLARITY (Contextual Linguistic Adaptation and Retrieval for Inclusive TTS sYnthesis), a backbone-agnostic framework that addresses these biases through dual-signal optimization: (i) contextual linguistic adaptation that localizes input text to the target dialect, and (ii) retrieval-augmented accent prompting (RAAP) that supplies accent-consistent speech prompts. Across twelve English accents, CLARITY improves accent accuracy and fairness while maintaining strong perceptual quality [^1] Code and audio samples are available at https://github.com/ICT-SIT/CLARITY.
Index Terms—  Instruction TTS, Accent Bias, Retrieval
1 Introduction
Instruction-guided TTS systems [du2024cosyvoice, lyth2024natural, zhou2024voxinstruct, wang2025spark] taking two inputs: 1) transcript text and 2) speaker descriptions of desired speech characteristics, have achieved promising instruction-following abilities. However, when prompted with natural language instructions, such TTS systems are susceptible to various biases. While gender bias has been examined by prior work [kuan-lee-2025-gender], to the best of our knowledge, no study has jointly addressed accent and linguistic bias. The two are interrelated but accent bias stems from the TTS system, while linguistic bias is often implicit in user input. The dual bias, as illustrated in Fig. 1, reduces generated speech authenticity by misaligning what is said and how it is spoken [burns2019speaking]. For example, instructing TTS to speak in a British Accent but with American English oriented text [^2] Note that traditional TTS assumes fixed text. This work addresses scenarios where users want to generate speech in a specific accent but cannot provide text in that variant of English. Hence, the input text must be slightly adapted to the accent specified in the speaker instruction..
Fig. 1: The “dual-bias” linguistic mismatch (prompt-source) with accent bias (system-source) reduces naturalness and authenticity.
Linguistic Bias – in the context of TTS, this is generally not an issue of model failure but a human-centric problem rooted in users’ ingrained linguistic knowledge, shaped by education and culture [clark2021impact]. It often leads to standardized input text, even when dialectal diversity is intended, preventing authentic synthesis. Recent studies confirm this tendency: for instance, ChatGPT frequently defaults to “standard” English when prompted with non-standard dialect text [fleisig-etal-2024-linguistic].
Accent Bias – arises from training data imbalance, where mainstream accents (e.g., American English) dominate and less common ones are under-represented. As a result, models may default to generic accents even when instructed otherwise. Recent research confirms performance disparities across accents in synthetic speech. For example, [10.1145/3715275.3732018] found significant quality gaps when cloning voices with different English accents, potentially reinforcing accent-based discrimination in AI speech. Broader evaluations like AudioTrust [li2025audiotrust] revealed that accent is a stronger source of unfair bias in audio models than age or gender, Current models generally default to American-accented speech profiles.
CLARITY: To jointly mitigate this dual bias, we present Contextual Linguistic Adaptation and Retrieval for Inclusive TTS sYnthesis (CLARITY). Our proposed CLARITY is a backbone-agnostic framework to enhance zero-shot TTS by providing richer guidance for both linguistic content and accent style. Its novelty lies in a two-signal optimization to jointly mitigate the two biases through: (i) an LLM-guided text adaptation module to localize input to the target dialect, and (ii) a metadata-driven retrieval mechanism to supply accent-consistent prompts.
Our Contributions. We introduce a sociolinguistic innovation in TTS by using large language models (LLMs) for on-the-fly text style transfer, leveraging their knowledge of dialects and slang to enhance authenticity. By pairing this with retrieval-augmented prompting, CLARITY achieves accent control and dialect-aware synthesis. To our knowledge, this is the first framework to jointly address user-side linguistic bias and system-side accent bias. Objective and subjective evaluations show that CLARITY improves fairness, accent fidelity, and overall authenticity.
2 The Proposed CLARITY
The CLARITY framework promotes inclusive TTS by mitigating linguistic and accent biases through two signals: an adapted text $x$ and an accent-consistent prompt $s$, as shown in Fig. 2. We cast this as a two-signal optimization problem, aiming to maximize both linguistic and accent fidelity under the synthesis constraint of a backbone TTS model. Formally, the objective is defined as
$\max_{x^{*},s^{*}}\;\;J_{\text{LLM}}(x^{*},m)\;+\;C(s^{*},m),$ (1)
$\quad\text{s.t.}\;\;\hat{y}=g_{\text{TTS}}(x^{*},s^{*}),$
where $J_{\text{LLM}}(x^{*},m)$ denotes the judgment score assigned by an LLM-as-a-Judge evaluating whether the adapted text $x^{*}$ aligns with the requested metadata $m$, and $C(s^{*},m)$ is the accent recognition confidence score that the retrieved prompt $s^{*}$ matches the target accent attributes. The synthesis function $g_{\text{TTS}}$ represents any zero-shot backbone TTS model.
LLM-Guided Instruction Parsing: Let $u\in\mathcal{U}$ be a free-form user instruction in natural language, and $m\in\mathcal{M}$ the structured metadata schema, $m=(\text{accent}\in\mathcal{A},;\text{gender}\in{M,F},;\text{age}\in\mathbb{N},\ldots)$, where $\mathcal{A}$ is the set of supported accents. The challenge is that $u$ may encode attributes explicitly, vaguely, or implicitly; for example, instead of the prescriptive instructions shown in Fig. 2, there may be one open to interpretation such as u=u=Read this like a local in Singapore” (a locale with multiple races and accents).
We employ an LLM to parse $u$ into the structured metadata $m$, with the following parsing function:
$m=f_{\text{parse}}(u;\theta_{\text{LLM}}),$ (2)
where $f_{\text{parse}}$ is realized by prompting the LLM with schema-specific instructions, and $\theta_{\text{LLM}}$ are its parameters. For each attribute $m_{j}$ (e.g., accent, gender, age), the LLM estimates a posterior distribution, $P(m_{j}\mid u;\theta_{\text{LLM}}),\quad m_{j}\in\mathcal{V}_{j},$ where $\mathcal{V}_{j}$ denotes the vocabulary/domain of attribute $j$. The predicted value is chosen as $m_{j}=\arg\max_{v\in\mathcal{V}_{j}}P(m_{j}=v\mid u;\theta_{\text{LLM}}).$ In cases where $u$ provides insufficient information to infer slot $m_{j}$, the system falls back to default $\pi_{j}$.
Contextual Linguistic Text Adaptation: Let $x$ denote the input text provided by the user, linguistic bias on the user side arises because $x$ is often drawn from the distribution of the user’s learned dialect $\mathcal{D}_{u}$, which may differ from the target dialect $\mathcal{D}_{t}$. We formalize text adaptation as a conditional function as follows:
$f_{\text{adapt}}:(x,m)\mapsto x^{\prime},$ (3)
where $f_{\text{adapt}}$ transforms the input $x$ into a culturally localized form $x^{\prime}$ that better reflects the target dialect $\mathcal{D}_{t}$.
To ensure that the adapted text $x^{\prime}$ is linguistically faithful to the target dialect $\mathcal{D}_{t}$, we generate candidate adaptations $\{x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{K}\}$ using different LLMs (e.g., GPT and LLaMA). We compare three candidates: the original standard text $x_{\text{std}}$ and adapted texts $k\in\mathcal{K}=\{\text{GPT},\text{LLaMA}\}$ produced by different LLMs. An LLM-as-a-Judge assigns each candidate a score $J_{\text{LLM}}(x,m)\in[0,10],$ which reflects how well $x$ aligns with the metadata $m$, particularly the accent attribute. The final text is then chosen as the candidate with the highest judgment score:
$x^{*}=\operatorname*{arg\,max}_{x\in\{x_{\text{std}}\}\,\cup\,\{x^{\prime}_{k}\}_{k\in\mathcal{K}}}J_{\text{LLM}}(x,m).$ (4)
Fig. 2: Our proposed CLARITY for mitigating user-side linguistic bias and system-side accent bias.
Retrieval-Augmented Accent Prompting (RAAP): System-side accent bias arises when a TTS backbone trained on imbalanced data defaults to majority accents $\mathcal{A}_{\text{dom}}$. To counter this, we introduce Retrieval-Augmented Accent Prompting (RAAP), which selects accent-consistent prompts from a curated pool $\mathcal{P}=\{(a_{i},z_{i},s_{i})\}$ containing metadata, transcript, and speech. Given target metadata $m$, we filter $\mathcal{P}$ by $a_{i}\approx m$ to form $\mathcal{P}_{m}$, then score each candidate $s_{i}$ using an ECAPA-TDNN [desplanques2020ecapa] for accent confidence $C(s_{i},m)\in[0,1]$, and select those with the highest scores to guide accented speech generation. Mathematically,
$s^{*}=\operatorname*{arg\,max}_{s_{i}\in\mathcal{P}_{m}}C(s_{i},m).$ (5)
To further align the text content of the prompt speech with the user’s text input, RAAP calculates the text similarity between the prompt speech candidates and the user-provided (standard) text input using $\cos(\phi(z_{i}),\phi(x))$, where $\phi(\cdot)$ is a TF–IDF [salton1988term] text embedding representation and $\cos(\cdot,\cdot)$ denotes cosine similarity. Specifically, we define the effective ranking score as
$r_{i}=C(s_{i},m)+\cos(\phi(z_{i}),\phi(x)),$ (6)
Backbone-agnostic Guided Synthesis: The final stage of CLARITY aligns the two signals: the adapted text $x$ and the accent-consistent prompt $s$, for guided synthesis. Let $g_{\text{TTS}}$ be the backbone TTS with parameters $\theta_{\text{TTS}}$. As a backbone-agnostic framework, CLARITY can use any instruction-guided, zero-shot TTS model; here we adopt the SOTA CosyVoice2 [du2024cosyvoice2] with its zero-shot functionality. The synthesis is defined as
$\hat{y}=g_{\text{TTS}}(x^{*},s^{*};\theta_{\text{TTS}}),$ (7)
where $g_{\text{TTS}}$ is steered through the two signals $x^{*}$ and $s^{*}$, generating final output $\hat{y}$ that aligns with both the localized text style and the accent-consistent prior.
3 Experiments
3.1 Experimental Settings
Accent Pool Construction: We evaluated our method on twelve English accents: ten accents (Canadian (CA), Chinese (CN), Spanish (ES), British (GB), Indian (IN), Japanese (JP), Korean (KR), Portuguese (PT), Russian (RU), and American (US)) are from the AESRC dataset (52,614 utterances from 528 speakers) [shi2021accented], each contributing 44–92 speakers and 4.2k–7k utterances, balanced by gender and ages 15–69. The remaining two accents are drawn from the SEAME dataset [lyu2010seame]: 40 Malaysian speakers (4,358 utterances, aged 20 to 33) and 114 Singaporean speakers (17,750 utterances, aged 18 to 24), with code-switching between English and Mandarin.
Text and Free-form Instructions Curation: For transcript, GPT-4333https://openai.com/index/gpt-4-research/ generated standard sentences across four scenarios (restaurant, university, workplace, supermarket). Speaker instructions derived from accent-pool metadata (accent, age, gender) served both as prompts and ground-truth labels for retrieval. Each instruction–text pair was then adapted by GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ and LLaMA-3.1-8B [^5] https://huggingface.co/meta-llama/Llama-3.1-8B to produce accent-specific variants. Fidelity was evaluated by GPT-5666https://openai.com/index/introducing-gpt-5/, which rated text–accent alignment on a 1–10 scale.
Baseline Systems: Two open-source SOTA prompt-based TTS systems are selected as baselines: ParlerTTS [lacombe-etal-2024-parler-tts, lyth2024natural] and CosyVoice2 [du2024cosyvoice]. Both take user instructions and standard text as input, but CosyVoice2 requires additional prompt speech even for instruction-guided functionality [^7] https://github.com/Render-AI/CosyVoice2 where a silent speech sample is provided.
Object Evaluation Metrics: To quantify the accent level of the generated speech, we calculate accent accuracy using the ECAPA-TDNN accent recognition model [^8] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa, finetuned on the accent pool (Section 3.1), as in RAAP. Speech quality is measured with NISQA [mittag2021nisqa] (1–5 scale) using its pre-trained model [^9] https://github.com/gabrielmittag/NISQA. Bias/Fairness of the generated speech is evaluated with the Fairness Discrepancy Rate (FDR) [de2021fairness, estevez2023study], which quantifies disparities in false alarm/reject rates across accents, with 1 indicating perfect fairness.
Subject Evaluation Metrics: We conducted human listening tests on four accented English samples (CN, SG, IN, GB). For each accent, 10 utterances (5F/5M) were generated from three systems: CosyVoice2 (standard text), ParlerTTS (standard text), and CLARITY. Native speakers of each accent (fluent in English), rated corresponding samples (e.g., Chinese listeners for Chinese-accented English [^10] The setting assumes that Chinese listeners are more sensitive to Chinese-accented English speech.). There were 4F/4M listeners for CN/SG and 2F/2M for IN/GB. Samples were evaluated on a 1–5 scale for naturalness, accent accuracy, age, gender consistency, and overall instruction match.
Fig. 3: RAAP accuracy (%) for gender, predicted accent, age attribute (left) and LLM-as-judge scores for various texts (right).
Fig. 4: Accent accuracy (%) (left) and NISQA (right) for the baselines and proposed CLARITY with GPT adapted text.
Table 1: Ablation study: RAAP with text similarity, accent confidence, and text adaptations. ‘Max’ means adapted text selection described in Eqn. 4.
Text Sim. Accent Score Text Apt. ACC (%) NISQA
45.60 4.21
46.26 4.28
61.74 4.25
LLaMA 62.47 4.33
GPT 63.36 4.31
Max 62.38 4.29
Fig. 5: Subjective Evaluation Results.
3.2 Experimental Results and Bias Analysis
Objective Evaluation:
1. RAAP and Contextual Linguistic Text Adaptation: Fig. 3 shows the RAAP accuracy (left) and the adapted scores generated by GPT-5 and LLaMA models (right). RAAP achieves nearly 100% accuracy for most accents in terms of accent and gender retrieval, while age prediction is less accurate but acceptable, given its known difficulty. For the LLM-as-judge scores, we observe that the standard text already achieves reasonable accent scores, with all scores above 6. After adaptation using GPT-4o-mini, four accents (GB, MY, SG, and US) show improved adaptation scores. Interestingly, these accents correspond to locations where the language of daily communciation tends to be English. A possible explanation is that for these speakers, the model had higher confidence in modifying the text, while the training corpus distribution likely reinforces this behaviour, which GPT-4 effectively captures. In contrast, LLaMA shows decreased accent scores after adaptation, although we doubt these results since the judge is GPT-5, not LLaMA, and continue to evaluate both adapted texts.
2. Comparision of Baselines and CLARITY: Fig. 4 plots accent accuracy and NISQA scores for generated speech across various systems, along with the averaged scores. CosyVoice2 and ParlerTTS achieve very low accent accuracy (8.75% and 6.14%), likely due to training sets lacking accented speech. Per-accent results reveal higher scores for accents better represented in training data. CLARITY with GPT-adapted text achieves the best overall accuracy, though ES, JP, and PT remain below 50%, showing the difficulty of these accents. In terms of NISQA, ParlerTTS averages 4.67, while CLARITY_GPT achieves 4.36, outperforming CosyVoice2 (3.16) and even ground truth on some accents.
3. Ablation Study of CLARITY: Table 1 presents the ablation study, showing that accent-score–guided prompt selection is highly effective, while adding text adaptation further improves accent accuracy. All variants achieve NISQA above 4.21, rising to 4.31 with RAAP and text adaptation, confirming the robustness of CLARITY. Among CLARITY variants, using GPT-adapted text performs best, while with LLaMA-adapted text is competitive, though its adapted text scores are lower (Fig. 3), indicating potential mismatches from using the GPT-5 as judge. Comparison of LLaMA and GPT for text selection does not improve results, suggesting that GPT-5 scores may not be the optimum [^11] Using multiple LLM-as-Judge models, e.g., GPT-5 and LLaMA-4, may reduce such mismatches, we leave this for future work.. Since GPT-adapted CLARITY performs best in the objective test, we adopt this setting for CLARITY in the following experiments.
Fig. 6: FDR comparison.
Fig. 7: Predicted accent distributions for baseline vs. proposed systems. TP = True Positive, FP = False Positive; correctly predicted accents are shadowed.
Subjective Evaluation: Fig. 5 shows violin plots of the listening test results. For naturalness, ParlerTTS performs best on IN, while CLARITY outperforms both alternatives on CN and GB. SG results are similar across systems, with CosyVoice2 slightly worse. This aligns with the NISQA scores in Fig. 4. CLARITY demonstrates superior accent consistency with reduced bias across all accents. Both CLARITY and ParlerTTS maintain gender consistency, whereas CosyVoice2 mimics prompt speech rather than following user instructions. Overall, CLARITY best aligns with user instructions.
Bias Analysis: The FDR across all accent groups in Fig. 6 shows that CLARITY systems with higher FDRs achieve greater fairness than the baselines. The distribution of true and false predictions across accents for CosyVoice2 and CLARITY in Fig. 7 indicates that CosyVoice2 (left) heavily biases toward US and CA accents (large dark-green and dark-red bars), often predicting them regardless of the real/true accent. CLARITY reduces this bias (right), decreasing US/CA dominance and increasing correctly predicted segments (shadowed), especially for GB and IN, resulting in more accurate and balanced accent distributions. This is particularly evident for GB and IN, where correctly predicted segments are much larger in CLARITY than in the baseline. Moreover, CLARITY achieves better subjective results in Fig. 5 for accent consistency and overall match scores, further confirming the effectiveness of the proposed method in mitigating accent bias.
4 Conclusion
The proposed CLARITY framework demonstrates that jointly mitigating accent and linguistic bias in instruction-guided TTS leads to measurable gains in inclusivity and authenticity. Across twelve English accents, CLARITY achieves the highest accent accuracy, competitive naturalness, and significantly improved fairness compared to state-of-the-art baselines. Human listening tests confirm reduced bias perception and stronger alignment with target accents. Although validated on twelve accents, the framework is inherently backbone-agnostic and can be easily extended to a wider range of accents, with future work focusing on multilingual adaptation and deeper accent–linguistic alignment.
Appendix A Appendix
In this section, we provide additional information about the proposed system, including details of the experimental protocol, as well as supplementary results on the accent pool data and generated speech.
A.1 Experimental Protocol
A.1.1 Accent Pool Data Generation
Table 2: The statistics of Accent Pool (AESRC)
Accent #Spk #Utt #F #M Age (min/max) CA 44 4,400 22 22 18/49 CN 50 5,000 26 24 17/38 ES 44 4,400 22 22 19/45 GB 92 9,031 42 50 16/65 IN 42 4,200 22 20 15/37 JP 46 4,600 23 23 18/69 KR 46 4,600 23 23 19/39 PT 53 5,283 27 26 18/62 RU 41 4,100 21 20 18/41 US 70 7,000 37 33 15/63
Table 3: The statistics of Accent Pool (SEAME)
Accent #Spk #Utt #F #M Age
(min/max)
MY 40 4,358 21 19 20/33
SG 114 17,750 63 51 18/24
Table 4: Overview of scenarios and metadata coverage in the curated instruction dataset.
Scenario Description Unique Metadata Combinations Approx. Instructions
Restaurant / Coffee Shop Informal, service-oriented interactions such as ordering or small talk. 575 $\sim$1,200
University / School Formal and instructional contexts such as lectures or classroom exchanges. 585 $\sim$1,200
Workplace / Office Professional settings emphasizing collaboration and presentation. 578 $\sim$1,200
Total 3,600
Fig. 8: Distributions of gender (left) and age (right) across all generated implicit instructions.
We evaluate our proposed method on twelve English accents. The accent pool is composed of accents from two datasets. Tables 2 and 3 summarize the statistics of the accent pool, including the number of speakers, utterances, gender distribution, and age range. Specifically, ten accents are drawn from the AESRC dataset, which was introduced in the Interspeech 2020 Accented English Speech Recognition Challenge [shi2021accented]. This dataset comprises approximately 180 hours of speech spanning the ten English accents. Each speaker reads sentences covering common conversational topics as well as human–computer interaction commands. For our experiments, we selected a subset of 52,614 utterances, balanced by gender and covering a broad age range. For each speaker, up to 100 longest utterances were retained to ensure consistent recording quality and sufficient duration per sample. The selected subset includes the following accents: Canadian (CA), Chinese (CN), Spanish (ES), British (GB), Indian (IN), Japanese (JP), Korean (KR), Portuguese (PT), Russian (RU), and American (US).
The remaining two accents, Malaysian (MY) and Singaporean (SG), are represented by 22,108 utterances selected from the SEAME dataset [lyu2010seame], a large-scale, code-switching English–Mandarin corpus collected from spontaneous conversations in Malaysia and Singapore. It contains speech from both Malaysian and Singaporean speakers, with detailed accent and speaker annotations. To construct our subset, we extracted English-dominant utterances from both conversational and interview recordings, retaining only those labeled as English or code-switched (EN and CS) in the transcript metadata. Purely Mandarin (ZH) segments were excluded. Each utterance was required to contain at least 5 words, with a maximum of 100 utterances per speaker, prioritizing longer segments. After extraction, only speakers with English accent confidence scores greater than 0.9 were retained. These filtering criteria yielded a balanced selection of Malaysian and Singaporean speakers across both genders and an age range of 18 to 33 years.
A.1.2 Instruction Curation
To develop realistic and context-sensitive guidance for instruction-driven speech synthesis, we curated a dataset of 3,600 implicit instructions that describe how a sentence should be spoken under different social and accent conditions. Each instruction captures paralinguistic elements such as tone, pacing, rhythm, and articulation, which are crucial for generating speech that aligns with human accent perception.
The dataset spans three conversational scenarios, each chosen to represent a distinct communicative domain where accent and prosody vary naturally. For each scenario, we defined ten standard sentences and paired them with twelve target accents (CA, CN, ES, GB, IN, JP, KR, MY, PT, RU, SG, US). Each accent was then combined with ten randomly sampled metadata combinations (age, gender, and language background), resulting in approximately 1,200 implicit instructions per scenario and 3,600 in total.
Table 4 summarizes the three scenarios and their corresponding metadata diversity. All implicit instructions were generated automatically using the GPT-4 language model [openai2023gpt4]. For each pair of a base sentence and sampled metadata, the model was prompted to provide paralinguistic guidance that specifies how the sentence should be spoken. The generation prompt was structured as follows:
Prompt Template: You are an expert linguist and voice coach. Given the sentence and the speaker’s accent, age, gender, and language background, describe how the speaker should deliver the sentence so that it sounds natural for that accent. Focus on tone, pacing, rhythm, and articulation rather than factual traits. Respond in one or two short sentences.
Generated implicit instruction example: Input Text:Can I have a cup of coffee, please?”
Metadata: {accent: SG, gender: female, age: 25}
Generated Instruction:Speak briskly and warmly, with a cheerful tone and rhythmic cadence typical of Singaporean English.
Each implicit instruction acts as a behavioral abstraction derived from the metadata. It does not restate all demographic attributes explicitly; instead, GPT-4 selectively emphasizes the subset that influences speech prosody and delivery.
Across all 3,600 instructions, the gender and age distributions are balanced, ensuring broad representation of speaker profiles for instruction-guided text-to-speech synthesis. In total, the dataset includes 1,804 male speakers (50.1%) and 1,796 female speakers (49.9%). The age range spans from 15 to 69 years, with a mean of 32.3 years, a median of 30.0 years, and a standard deviation of 11.9. Figure 8 illustrates the gender and age distributions within the curated dataset.
Fig. 9: Pipeline from user instructions to LLM-as-a-judge
A.1.3 Adapted Text Generation
To parse user instructions, generate localised text, and select the best final spoken text, a total of four LLMs were used. Figure 9 illustrates the entire process from the initial user instruction to the final selection of the localised text.
Gemini-2.5-flash-lite was prompted to parse the user’s instructions into a structured representation. An example of the prompt used is shown below:
Example prompt used for instruction parsing You are given a natural language input from a user describing a desired speech generation request. Your job is to extract structured metadata suitable for retrieving a matching voice sample from a database.
User Instructions: {user_instructions}
User Text: {user_text}
The available accents in the dataset pool are: CA, CN, ES, GB, IND, JP, KR, PT, RU, US, SG, MY.
Please extract the following fields from the input. If not explicitly stated, make a reasonable inference based on the speech to be spoken. Otherwise, you may mark it as ”unspecified”.
Format your output as a JSON object like this: { ”accent”: ”Speaker accent. Choose from the above pool. If there is no exact match, choose the closest available option.”, ”language”: ”language in the phrase to be spoken, e.g. EN”, ”age”: ”exact_age — [range_start, range_end] – use exact age if given, e.g. 25; if approximate age only, infer a 10-year range, e.g. [20, 30].”, ”gender”: ”M — F”, ”tone”: ”e.g. soft, angry, romantic, etc.”, ”emotion”: ”optional emotion, e.g. love, sadness, happiness”, ”additional_context”: ”any inferred intent or style, e.g. persuasive, affectionate, instructional” }
Next, GPT-4o-mini and Llama-3.1-8B utilised the parsed speaker profile to perform text localisation independently. Each model produced a distinct version of localised text for subsequent evaluation and TTS generation. The prompt was refined iteratively to balance localisation accuracy with the downstream TTS processing constraints. The main goal was to ensure constraints such as ASCII-only output, preventing translation and ensuring structured output, while encouraging the LLM to naturally integrate local expressions into the text. The final version of the LLM prompt is shown below:
Example prompt used for text adaptation You are a text localiser, and your goal is to produce localised, accented speech meant for a TTS LLM to use. Speaker Information: {structured_instructions} User text to be spoken: {user_text} Instructions:
- DO NOT translate the text to any other language. Always keep it in the same language (e.g., English).
- Strictly use only ASCII characters in your adapted text.
- You may alter the sentence to fit how the local speaker would use the given language.
- You may add local expressions or discourse particles ONLY if it flows naturally in the local language.
- Ensure that expressions added are suitable with the tone and context of the sentence.
Return a JSON object with only the localized ”text” field: {{ ”text”: ”<<localized text here>>.” }}
Finally, GPT-5-mini was employed as an LLM-as-a-Judge, evaluating the standard text and both adapted versions. It provided quantitative scores and qualitative reasoning for each output. An example of the judging prompt is shown below:
Example prompt used for LLM-as-a-judge scoring You are an expert linguist and speech evaluator. Your task is to score how well each given text sample is appropriate for a speaker from a specific accent or region. Speaker info:
{speaker_info_structured}
Samples to evaluate:
{samples_list_str}
From the above, please derive the following:
- ”score”: an integer from 0 to 10, where 0 = completely inappropriate localisation for this speaker and 10 = appropriate and excellent regional adaptation with authentic local vocabulary, expressions, or language patterns.
- ”reason”: a short explanation of why the score was given, considering whether the language is natural for this speaker and noting any regional vocabulary or expressions.
IMPORTANT: Reward localisation only when it feels natural and contextually appropriate. The presence of a local particle or slang word does not automatically make a sentence better. If the expression feels forced, awkward, or out of place, the score should be reduced accordingly. Format your output as a JSON object like this: {{ ”score”: <<a list of scores in order of the samples>>, ”reason”: ”<<a list of short reasonings in order of the samples>>” }}
Fig. 10: Combined visualization of per-accent accuracy (left) and confusion matrix (right) on the validation set.
Fig. 11: Average NISQA for each accent category. Higher scores indicate better perceptual quality.
A.2 Additional Results on Accent Pool data
A.2.1 Accent Recognition Accuracy
To evaluate the performance of the accent preservation on the generated speech, we fine-tuned ECAPA-TDNN model, pretrained on the CommonAccent corpus [zuluaga2023commonaccent], for 12-class accent classification (CA, CN, GB, IN, JP, KR, MY, PT, RU, SG, ES, US) using the AESRC2020 and SEAME datasets. The model extracts 80-dimensional log-Mel filterbank features (25 ms window, 10 ms hop) followed by mean-variance normalization. The ECAPA-TDNN backbone (embedding dimension = 192, $\sim$6.2 M parameters) was kept frozen, and a single linear layer with 12 output units was fine-tuned. We used a batch size of 8, AdamW optimizer with a learning rate of $1\times 10^{-4}$, and AAM loss as the objective. The model was trained for 15 epochs with a random seed of 42 on an NVIDIA RTX 4070 GPU. The dataset was randomly divided into 80% for training and 20% for validation.
Figure 10 presents both the per-accent classification accuracy and the confusion matrix on the validation set of the accent pool data. The model achieves high recognition performance for several distinct accent categories, such as Indian (IN, 99%), Singaporean (SG, 91%), and British (GB, 89%) English. However, recognition accuracy decreases noticeably for Japanese (JP, 50%), Chinese (CN, 59%), and Portuguese (PT, 57%) accents, reflecting the challenges posed by their acoustic and phonological proximity to other varieties.
The confusion patterns reveal regional clustering effects. East Asian accents (JP, CN, KR) exhibit substantial inter-class confusion due to shared phonetic and prosodic characteristics, such as vowel centralization and simplified consonant clusters. North American accents (CA, US) show bidirectional confusion, which aligns with their linguistic proximity and overlapping vowel inventories. Similarly, European accents (ES, PT, RU) demonstrate cross-lingual confusion patterns that may arise from similar rhythmic and prosodic influences across Romance and Slavic language backgrounds.
These findings indicate that the accent recognition model performs well across most accents and provides reliable judgments on both original and generated accent speech.
A.2.2 Speech Quality Evaluation
Figure 11 shows the average NISQA score for each accent on the accent pool data. The generated speech maintains high perceptual quality across all accents, with scores ranging from 4.0 to 4.6 on a five-point scale. Singaporean (SG, 4.57), Japanese (JP, 4.41), and British (GBR, 4.40) accents achieve the best quality, indicating clear and natural synthesis. Spanish (ESP, 4.00) and Chinese (CHN, 4.13) exhibit slightly lower scores, likely due to accent-specific variability or limited training coverage. Despite these differences, the narrow score range demonstrates strong cross-accent generalization and consistent perceptual quality.
These findings suggest that the accent pool data is of sufficient quality to be used as prompt speech in the speech generation stage.
Table 5: Ablation study of RAAP, comparing text similarity using standard user-provided text or LLM-adapted text.
Text Sim. Accent Score Text Apt. ACC (%) NISQA
45.60 4.21
Adapted LLaMA 62.23 4.30
Standard LLaMA 62.47 4.33
Adapted GPT 62.26 4.35
Standard GPT 63.36 4.31
Fig. 12: Subjective evaluation results for GPT-based methods using either adapted or standard text for text similarity in prompt speech selection.
Fig. 13: Female evaluators
Fig. 14: Male evaluators
Fig. 15: Predicted accent distributions for baseline vs. proposed systems. TP = True Positive, FP = False Positive; correctly predicted accents are shadowed.
A.3 Additional Results on Generated Speech
A.3.1 Results on Retrieved Prompt Speech
To align the text content of the prompt speech with the user’s input, RAAP calculates the text similarity between the prompt speech candidates and the user-provided (standard) text input, as explained in Eqn. 6. Alternatively, the adapted text can also be used as the query to help identify prompt speech with similar content.
The objective results are presented in Table 5, where adapted text yields slightly lower accent accuracy for both LLaMA and GPT, while achieving comparable NISQA scores. In the subjective listening test, GPT-based methods using either adapted or standard text were evaluated. The violin plots in Figure 12 show that for SG, the results are almost identical between standard and adapted text; for CN, a similar trend is observed except that MOS with standard text achieves slightly better performance; while for GB and IN, GPT with standard text as the basis for text similarity calculation tends to obtain higher scores across all perceptual evaluation metrics.
A.3.2 Subjective Analysis
Figure 14 shows the subjective results for female (top) and male (bottom) evaluators separately. The female results exhibit slim violins, indicating high agreement among listeners as ratings are clustered, whereas the male results show wider (fatter) violins, indicating lower agreement and more dispersed ratings, particularly for CN and SG accent speech.
A.3.3 Bias Analysis
Statistical Analysis - Binomial Test: to specifically test for the baseline model’s bias toward US and CA accents, we use a Binomial test. The event in this context is: a true non-US/CA accent sample being incorrectly predicted as a US/CA accent by the model. Our Null hypothesis (H_0H\_0) is as follows: The probability of the model predicting a non-US/CA accent as a US/CA accent is equal to the random chance probability (with 12 total accents, the random chance is $1/12\approx 0.083$). This represents no model bias. On the other hand, if the probability of the model predicting a non-US/CA accent as a US/CA accent is significantly higher than the random chance probability, it indicates that the model has a bias.
Table 6: Binomial test: biases towards US and CA accents comparision for CosyVoice2 and CLARITY
Accent CosyVoice2 CLARITY_GPT P-value_US P-value_CA P-value_US P-value_CA CN 3.60E-67 2.25E-05 6.86E-4 0.45 ES 2.93E-79 1.10E-08 0.01 0.36 GB 2.80E-64 7.49E-11 1 1 IN 2.65E-80 2.12E-10 1 1 JP 4.03E-74 9.51E-05 0.29 0.29 KR 2.65E-80 9.52E-07 0.87 1.18E-3 MY 2.38E-114 4.68E-05 0.94 1 PT 1.53E-84 1.59E-09 8.14E-3 0.49 RU 3.69E-76 1.06E-05 0.01 1 SG 8.86E-88 1.88E-04 0.87 0.87
The P-value  [^12] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.binomtest.html of US and CA accents are shown in Table 6. The P-value indicates the result of the Binomial test, where a very small P-value (typically $p\leq 0.05$) means the observed proportion is significantly higher than the random chance, thus confirming the existence of a statistically significant bias. We can see that CosyVoice2 has a extreme and consistent bias towards US/CA accents across all tested non-US/CA accents. CLARITY_GPT demonstrates a much lower and more variable bias compared to CosyVoice2 which reduced overall bias. For British (GB), India (IN), Japan (JP), Malaysian (MY) and Singapore (SG) accents, there is no statistically significant evidence of bias toward US/CA accents. Although the magnitude of the bias is lower, CLARITY_GPT still exhibits a statistically significant bias for several accents, particularly Chinese (CN), Spanish (ES), Korean (KR) and Portuguese (PT). Bias towards US accent is more significant than CA accent.
Ablation Study: as shown in Figure 15, this ablation study details the impact of various model configurations on accent accuracy and systemic bias, demonstrating a phased improvement across experimental setups. The initial phase, which involved replacing the near-silent prompt audio with accent pool sourced prompt audio dynamically selected via instruction extracted speaker features, alone yielded a significant boost in both accent accuracy and fairness metrics. Building upon this foundation, performance saw further incremental gains through the subsequent integration of target text (standard) to prompt transcription similarity and the strategic use of prompt speech exhibiting a more pronounced accent (higher accent score). As mentioned in Section A.3.1, we also compared the results of standard and adapted text similarity. Ultimately, the most effective configurations (proposed) were achieved by leveraging Llama or GPT to perform adaptation on the target text, resulting in the highest overall accuracy and the lowest observed degree of bias among all tested methodologies.